OpenStudy (anonymous):

5a+15b=5(a+3b) factorize.........hlp please

OpenStudy (anonymous):

you have already factorised it

OpenStudy (aroub):

5a+15b =5a+15b

OpenStudy (anonymous):

@aroub it is not factorisation it is solving

OpenStudy (sriram):

hashir is right

OpenStudy (aroub):

yea i know!! its is already factored i guess!! :)

OpenStudy (anonymous):

yup medal please

OpenStudy (anonymous):

thanxxxxxxxxx aroub >d but ya solved it not factorized ur right aroub its factorized

OpenStudy (anonymous):

5(a+3b)=5(a+3b)

OpenStudy (anonymous):

Now it's factored.

OpenStudy (sriram):

i guess do2 deserves a medal for this

OpenStudy (anonymous):

thnxxxxxx siriam >d>d>d